OpenStudy (anonymous):

Solve by elimination: 9x2 + y2 = 85 2x2 - 3y2 = 6

OpenStudy (anonymous):

Eliminate y by multipling both sides of the first equation by 3 and then add the equations together: 27x^2 +3y^2 = 255 2x^2 - 3y^2 = 6 -------------------- -29x^2 = 261 x^2 = 9 Knowing x you can now solve for y by plugging x into either equation.\[\{\{x\to -3,y\to -2\},\{x\to -3,y\to 2\},\{x\to 3,y\to -2\},\{x\to 3,y\to 2\}\} \]

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

My mistake. The negative sign of -29x^2 should not be there.

OpenStudy (anonymous):

Thank you for the medal.